OpenStudy (shadyykatie):

4. Find the missing value to the nearest hundredth. sin = _____ = 7/12 5. Find the missing value to the nearest hundredth. cos = ______ = 9/20

OpenStudy (anonymous):

sin(x) = 7/12 x = arcsin(7/12) x = ???